The Wrigley Corporation needs to raise $32 million. The investment banking firm of Tinkers, Evers & Chance will handle the transaction.


a. If stock is utilized, 1,800,000 shares will be sold to the public at $19.40 per share. The corporation will receive a net price of $18.00 per share. What is the percentage underwriting spread per share? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)
  


b. If bonds are utilized, slightly over 32,400 bonds will be sold to the public at $1,002 per bond. The corporation will receive a net price of $995 per bond. What is the percentage of underwriting spread per bond? (Relate the dollar spread to the public price.) (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)
  


c-1. Which alternative has the larger percentage of spread?
  

Stock
Bond


c-2. Is this the normal relationship between the two types of issues?
  

Yes
No
Answer #1
a) percentage of underwriting spread per bond = (19.4-18)/19.4 = 7.22%
b) percentage of underwriting spread per bond= (1002-995)/1002 =0.70%
c-1) stock has the larger percentage of spread
c-2) Yes